OpenStudy (zubhanwc3):

use riemann sum for (sinx) [0,2pi]

OpenStudy (zubhanwc3):

\[x_{0}=0/x _{1}=\pi/4/x _{2}=\pi/3/ x _{3}=\pi/ x _{4}=2\pi\]

OpenStudy (zubhanwc3):

\[c _{1}=\pi/6\]

OpenStudy (zubhanwc3):

\[c _{2}=\pi/3/c _{3}=2\pi/3/c _{4}=3\pi/2\]

OpenStudy (abb0t):

First, you need to find \(\sf \Delta x = \frac{b-a}{n}\), where you have: \(\huge \lim_{n \rightarrow \infty} \sum_{i=1}^{n} f(x_i)\Delta x\)

OpenStudy (zubhanwc3):

how do i find n?

OpenStudy (abb0t):

Use limits.